OpenStudy (anonymous):

Given the function f(x) = 2(x + 6), find x if f(x) = 22.

OpenStudy (anonymous):

2(x+6)=22 solve for x

OpenStudy (anonymous):

You are given with f(x).. Plug in the value in the equation: \[2(x+6) = 22\] \[x+6 = 11\] Subtract 6 both sides and find the values of x..

OpenStudy (anonymous):

but this is my answer choices −36 −7 52 −15

OpenStudy (lgbasallote):

hah @waterineyes made a mistake :p

OpenStudy (anonymous):

Are you 100% sure about the function? And the answer choices?

OpenStudy (anonymous):

yes im sure

OpenStudy (lgbasallote):

actually she mistakenly wrote 5 as 52 lol

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

never mind lol these r the choices 56 5 38 17

OpenStudy (anonymous):

5 is the correct one.

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Does that makes sense? Any doubt? Please, ask! :)

OpenStudy (lgbasallote):

darn.. @waterineyes was so close to being wrong :p

OpenStudy (anonymous):

lol wanna answer another one for me ?

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

If h(x) = −2x − 10, find h(−4). −2 −18 −3 −16

OpenStudy (anonymous):

h(-4)=-2*(-4)-10 Your turn

OpenStudy (anonymous):

Come on! You can ! Give it a try!

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

Excellent! Well done!

OpenStudy (anonymous):

Want to try any other? Please post it for the community! ;)

OpenStudy (anonymous):

huh wats that mean ?

OpenStudy (anonymous):

@lgbasallote good to see your comment in openstudy.. Long time it has been..